# MGMCpaper.bib

### v1

@@ Abstract
Gaussian random fields play an important role in many areas of science and engineering. In practice, they are often simulated by sampling from a high-dimensional multivariate normal distribution which arises from the discretisation of a suitable precision operator on a finite grid. Existing methods such as Cholesky factorisation and Gibbs-sampling become prohibitively expensive on fine meshes due to their high computational cost or the fact that they do not explore the probability space efficiently. In this work we revisit the Multigrid Monte Carlo algorithm of Goodman & Sokal (Physical Review D 40.6, 1989), which can overcome these issues. The novelty of our work consists in the application of the method to linear Bayesian settings where the prior is constrained by a finite set of observations. For this, we develop a bespoke random smoother which takes care of the low-rank updates that arise in Bayesian inference. We provide a rigorous analysis of the method based on the link between linear solvers and samplers for multivariate normal distributions, drawing on standard multigrid convergence theory in a finite element setting. In particular, we prove that Multigrid Monte Carlo is algorithmically near-optimal in the limit of the grid-size going to zero. These theoretical results are confirmed by numerical experiments which demonstrate that Multigrid Monte Carlo can be significantly more efficient than alternative methods when applied in a Bayesian setting.
@@ 1 Introduction
The approximate simulation of Gaussian random fields plays a pivotal role in a large number of research areas, such as quantum physics [goodman1989multigrid, edwards1992multi], spatial statistics [LindgrenEtAl.F_2011_ExplicitLinkGaussian], additive manufacturing [ZhangEtAl.H_2021_StochasticModelingGeometrical], cosmology [Marinucci.D_2011_book], natural language processing [inoue-etal-2022-infinite], public health [MannsethEtAl.J_2021_VariationUseCaesarean], geosciences [LiuEtAl.Y_2019_AdvancesGaussianRandom], or uncertainty quantification in engineering applications [Soize.C_2017_UncertaintyQuantificationAccelerated, VasileEtAl.M_2021_AdvancesUncertaintyQuantification]. In theory, it is a well understood and solved problem, but in practice the efficiency and performance of traditional sampling procedures degenerates quickly when the random field is discretised on a grid (lattice) with spatial resolution going to zero. Moreover, most existing algorithms do not scale well on large-scale parallel computers. Direct approaches based on factorisations of the covariance or precision matrix, e.g., [rue2001fast, rue2005gaussian], run into memory problems and their cost per sample scales in general like $\mathcal{O}(n^{1+\zeta})$ with respect to the number of grid points $n$ where $\zeta>0$ is a positive constant; the computational complexity can only be reduced to $\mathcal{O}(n\log(n))$ if the matrix is structured and FFT-based approaches are applicable. Additional complications arise in the case where the Gaussian distribution is conditioned on data, as in (linear) Bayesian inverse problems, or when the spectrum of the covariance operator decays slowly. On the other hand, stationary, iterative approaches, such as random-walk Metropolis-Hastings [metropolis1953, hastings1970], preconditoned Crank–Nicholson [cotter2013mcmc], Langevin-based samplers [roberts1996exponential] or Hamiltonian Monte Carlo [duane1987216] become extremely inefficient as the grid size goes to zero. In particular, the convergence rate and the integrated autocorrelation time of the resulting Markov chain degenerates, leading to poor effective sample sizes. A lot of work has gone into this problem, but it is still a topic of ongoing research [LindgrenEtAl.F_2011_ExplicitLinkGaussian, LiuEtAl.Y_2019_AdvancesGaussianRandom, LangEtAl.A_2011_FastSimulationGaussian, Croci.M_etal_2018_EfficientWhiteNoise, FeischlEtAl.M_2018_FastRandomField, BachmayrEtAl.M_2020_UnifiedAnalysisPeriodizationBased, BolinEtAl.D_2020_RationalSPDEApproach, KressnerEtAl.D_2020_CertifiedFastComputations, LindgrenEtAl.F_2022_SPDEApproachGaussian].
@@
One of the most basic stationary iterative methods is the Gibbs sampler, see e.g. [RobertEtAl.CP_2004_MonteCarloStatistical, GelmanEtAl.A_2013_BayesianDataAnalysis, RubinsteinEtAl.RY_2016_SimulationMonteCarlo]. The problem of this standard method is its slow convergence to the target distribution and the strong correlation of the resulting samples, both of which become substantially worse as the resolution increases. Already in the late 1980s, Goodman, Sokal and their collaborators wrote a series of papers aimed at accelerating the Gibbs sampler using multigrid ideas, leading to extensive research activity in the quantum physics community [goodman1989multigrid, edwards1992multi, edwards1991multi, mana1996dynamic, mendes1996multi]. The key observation in the seminal paper by Goodman and Sokal [goodman1989multigrid] is the connection of random samplers, such as the Gibbs method, to iterative solvers for linear systems. In particular, the Gibbs sampler for generating Gaussian samples $x\sim\mathcal{N}(\overline{x},A^{-1})$ with mean $\overline{x}$ and covariance matrix $A^{-1}$ can be written as an iteration identical, except for an additive noise component, to the Gauss–Seidel iteration for solving the linear system $A\overline{x}=f$ to obtain $\overline{x}=A^{-1}f$; for more details see [goodman1989multigrid], as well as [Fox.C_Parker_2017_AcceleratedGibbsSampling]. Based on this observation, they proposed the so-called Multigrid Monte Carlo (MGMC) method—a random analogue of the multigrid method for solving discretised partial differential equations (PDEs). A careful analysis of the MGMC method is the focus of this paper. We assume that we have a positive definite operator $\mathcal{A}:H\to H$ on some function space $H$ on $\mathbb{R}^{d}$, for example $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ with positive integer $\alpha$ where $\Delta$ is the Laplace operator. A sufficiently accurate discretisation of $\mathcal{A}$ on some grid in $\mathbb{R}^{d}$ results in a possibly very large matrix $A\in\mathbb{R}^{n\times n}$. The goal of this paper is to show how to efficiently generate samples from $\mathcal{N}(\overline{x},A^{-1})$ for $n\rightarrow\infty$. We provide a rigorous theoretical justification for the optimal scaling of our method for large scale problems: the cost for generating an independent sample grows linearly with the problem size and the sampler is near-optimal in the continuum limit. Our analysis includes the important setting of a Gaussian random field conditioned on noisy data via a Bayesian approach; here, $\mathcal{A}$ is a finite-rank perturbation of some differential operator and the quantity of interest a functional of the Gaussian random field. Somewhat surprisingly, MGMC has not yet been considered in this context.
@@
    An extension of the MGMC algorithm to the important situation of sampling Gaussian random fields conditioned on noisy data, i.e., to linear Bayesian inverse problems.
@@
    A grid-size-independent convergence theory for the MGMC method in this Bayesian setting. The analysis shows that the first two moments, which fully characterise the Gaussian distribution, converge exponentially at a uniform rate to the target moments.
@@
    A corollary on the near-optimality of the cost of the MGMC method to produce approximate samples of the limiting infinite-dimensional Gaussian random field with a prescribed tolerance on the discretisation error.
@@
    An extension of the MGMC algorithm, as well as its convergence theory, to the case of conditonal Gaussian random variables.
  6. 6.
@@
It should be stressed that in contrast to single-grid stationary iterative methods such as the Gibbs sampler, the convergence rate of MGMC does not degenerate as the grid size tends to zero. The convergence analysis follows Fox and Parker [Fox.C_Parker_2017_AcceleratedGibbsSampling], who demonstrated that the convergence of the standard component-sweep Gibbs sampler is (in a suitable sense) equivalent to the deterministic (lexicographical) Gauss–Seidel iterative method, and that any matrix splitting used to generate a deterministic relaxation scheme also induces a stochastic relaxation that is a generalised Gibbs sampler; see also [norton2016tuning] where this analysis is extended also to non-Gaussian targets and to the popular Langevin and hybrid (or Hamiltonian) Monte Carlo samplers. Here, we extend the analysis in [Fox.C_Parker_2017_AcceleratedGibbsSampling] to MGMC. In doing so, we need to go beyond the derivation in [goodman1989multigrid] and beyond the standard multigrid theory (as presented for example in [Hackbusch.W_2016_IterativeSolutionLarge]) which provides convergence of the mean. Extra work is necessary to obtain the convergence of the covariance and to extend the analysis to the conditional case. In particular, for linear functionals of the random field that are bounded in $H$, the crucial approximation and smoothing properties for the conditional case can be reduced to the corresponding properties in the unconditional case.
@@ 2.2 Linear Bayesian inverse problem
We are interested in generating samples from the posterior distribution $\mathcal{N}(\mu,\widetilde{\mathcal{A}}^{-1})$ and this paper focuses on the multigrid Monte Carlo sampler, which was introduced in [goodman1989multigrid] and is described in Section 3 below. The generated samples can be used to compute the expected value of a quantity of interest which is a functional $\mathcal{F}:H\rightarrow\mathbb{R}$. This could for example be a bounded linear functional of the form
@@
where $\chi$ is an element in $H$.
@@ 2.3 Discretisation
To realise such random variables, we follow [DaPrato.D_2014_book, Section 4.1.2] and let $(\varphi_{j})_{j\geq 1}$ be a complete $H$-orthonormal system (ONS) consisting of eigenfunctions of $\mathcal{A}^{-1}$ with corresponding eigenvalues $(1/\lambda_{j})_{j\geq 1}$. Then, $(\varphi_{j}/\sqrt{\lambda_{j}})_{j\geq 1}$ is a complete ONS of $V=D(\mathcal{A}^{1/2})\subset H$. For each $\chi\in H$, we define
@@
the series in (2.20) converges in $L^{2}(\Omega)$ and the resulting random variable $(v,\chi)$ has the desired mean and covariance in (2.18)–(2.19). Moreover, following again [DaPrato.D_2014_book] the random field $v$ can be directly constructed on a suitably large Hilbert space without taking the coupling with $\chi\in H$.
@@ 4 Theory
This section presents our main theoretical results. We first put the relationship between random samplers and linear solvers already mentioned in Section 3.3 into a wider context: in Section 4.1 we argue that according to the pioneering work in [Fox.C_Parker_2017_AcceleratedGibbsSampling], the convergence of the Multigrid Monte Carlo sampler is closely related to the convergence of deterministic multigrid solvers for linear systems of equations. This is made more explicit in Section 4.2, where we show that the Multigrid Monte Carlo update Alg. 1 leaves the target distribution invariant, and Section 4.3, where we prove grid-independent convergence of the algorithm. Throughout the discussion we highlight the connection with the theory for deterministic multigrid methods; to streamline the presentation we defer proofs to the appendices. Finally, in Section 4.4, we apply and extend the theory to conditional distributions and Bayesian inverse problems introduced in Section 2.2. We conclude with a result on the near-optimality of the MGMG sampler for drawing samples from the target distribution in Section 4.5.
@@ 4.4 Application to linear Bayesian inverse problems
$x^{\top}\widetilde{M}_{\ell}^{\text{(SGS)}}x$ $=x^{\top}\widetilde{A}_{\ell}x+x^{\top}L_{\ell}\left(D_{\ell}+B_{\ell}\Gamma^{-1}B_{\ell}^{\top}\right)^{-1}L_{\ell}^{\top}x\geq x^{\top}\widetilde{A}_{\ell}x+x^{\top}L_{\ell}D_{\ell}^{-1}L_{\ell}^{\top}x\geq x^{\top}\widetilde{A}_{\ell}x\,.$
Since the matrix $\widetilde{A}_{\ell}$ was assumed to be positive definite, we also have $\widetilde{A}_{\ell}>0$.
@@ 4.4.1 Approximation property for the perturbed problem
In the lemma above we assumed $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for the unpurterbed precision matrix. This is a standard assumption
@@ 4.5 Optimality of MGMC
The MGMC update in Alg. 1 is efficient in the sense that is specified in the below Corollary 4.21. For this we consider the Bayesian setting in Section 4.4 with a symmetric positive definite operator $\widetilde{\mathcal{A}}$ that has a compact inverse (which is not necessarily of trace class) and vectors $f_{L}$ and $F_{L}$ that are defined as in (2.10) and (2.16) respectively; we remind the reader that these two vectors are constructed with the help of the bounded linear operators $\mathcal{B}$ and $\mathcal{F}$ introduced in Section 2.2.
@@
Then, the Multigrid Monte Carlo algorithm is optimal in the limit $n_{L}\rightarrow\infty$ in the sense that:
@@
    Generating a new sample in the Markov chain incurs a cost that grows linearly in the number of unknowns $n_{L}$ as $n_{L}\to\infty$.

### v2

@@ Abstract
Gaussian random fields play an important role in many areas of science and engineering. In practice, they are often simulated by sampling from a high-dimensional multivariate normal distribution, which arises from the discretisation of a suitable precision operator. Existing methods such as Cholesky factorization and Gibbs sampling become prohibitively expensive on fine meshes due to their high computational cost. In this work, we revisit the Multigrid Monte Carlo (MGMC) algorithm developed by Goodman & Sokal (Physical Review D 40.6, 1989) in the quantum physics context. To show that MGMC can overcome these issues, we establish a grid-size-independent convergence theory based on the link between linear solvers and samplers for multivariate normal distributions, drawing on standard multigrid convergence theory. We then apply this theory to linear Bayesian inverse problems. This application is achieved by extending the standard multigrid theory to operators with a low-rank perturbation. Moreover, we develop a novel bespoke random smoother which takes care of the low-rank updates that arise in constructing posterior moments. In particular, we prove that Multigrid Monte Carlo is algorithmically optimal in the limit of the grid-size going to zero. Numerical results support our theory, demonstrating that Multigrid Monte Carlo can be significantly more efficient than alternative methods when applied in a Bayesian setting.
@@ 1 Introduction
The approximate simulation of Gaussian random fields plays a pivotal role in a large number of research areas, such as quantum physics [goodman1989multigrid, edwards1992multi], spatial statistics [LindgrenEtAl.F_2011_ExplicitLinkGaussian], additive manufacturing [ZhangEtAl.H_2021_StochasticModelingGeometrical], cosmology [Marinucci.D_2011_book], natural language processing [inoue-etal-2022-infinite], public health [MannsethEtAl.J_2021_VariationUseCaesarean], geosciences [LiuEtAl.Y_2019_AdvancesGaussianRandom], or uncertainty quantification in engineering applications [Soize.C_2017_UncertaintyQuantificationAccelerated, VasileEtAl.M_2021_AdvancesUncertaintyQuantification]. In theory, it is a well understood and solved problem, but in practice the efficiency and performance of traditional sampling procedures degenerates quickly when the random field is discretised on a grid (lattice) with spatial resolution going to zero. Moreover, most existing algorithms do not scale well on large-scale parallel computers. Direct approaches based on factorisations of the covariance or precision matrix, e.g., [rue2001fast, rue2005gaussian], run into memory problems and their cost per sample scales in general like $\mathcal{O}(n^{1+\zeta})$ with respect to the number of grid points $n$ where $\zeta>0$ is a positive constant; the computational complexity can only be reduced to $\mathcal{O}(n\log(n))$ if the matrix is structured and FFT-based approaches are applicable. Additional complications arise in the case where the Gaussian distribution is conditioned on data, as in (linear) Bayesian inverse problems, or when the spectrum of the covariance operator decays slowly. On the other hand, stationary, iterative approaches, such as random-walk Metropolis-Hastings [metropolis1953, hastings1970], preconditoned Crank–Nicholson [cotter2013mcmc], Langevin-based samplers [roberts1996exponential] or Hamiltonian Monte Carlo [duane1987216] become extremely inefficient as the grid size goes to zero. In particular, the convergence rate and the integrated autocorrelation time of the resulting Markov chain degenerate, leading to poor effective sample sizes. A lot of work has gone into this problem, but it is still a topic of ongoing research [LindgrenEtAl.F_2011_ExplicitLinkGaussian, LiuEtAl.Y_2019_AdvancesGaussianRandom, LangEtAl.A_2011_FastSimulationGaussian, Croci.M_etal_2018_EfficientWhiteNoise, FeischlEtAl.M_2018_FastRandomField, BachmayrEtAl.M_2020_UnifiedAnalysisPeriodizationBased, BolinEtAl.D_2020_RationalSPDEApproach, KressnerEtAl.D_2020_CertifiedFastComputations, LindgrenEtAl.F_2022_SPDEApproachGaussian].
@@
One of the most basic stationary iterative methods is the Gibbs sampler, see e.g. [RobertEtAl.CP_2004_MonteCarloStatistical, GelmanEtAl.A_2013_BayesianDataAnalysis, RubinsteinEtAl.RY_2016_SimulationMonteCarlo]. The problem of this standard method is its slow convergence to the target distribution and the strong correlation of the resulting samples, both of which become substantially worse as the resolution increases. Already in the late 1980s, Goodman, Sokal and their collaborators wrote a series of papers aimed at accelerating the Gibbs sampler using multigrid ideas, leading to extensive research activity in the quantum physics community [goodman1989multigrid, edwards1992multi, edwards1991multi, mana1996dynamic, mendes1996multi]. The key observation in the seminal paper by Goodman and Sokal [goodman1989multigrid] is the connection of random samplers, such as the Gibbs method, to iterative solvers for linear systems. In particular, the Gibbs sampler for generating Gaussian samples $x\sim\mathcal{N}(\overline{x},A^{-1})$ with mean $\overline{x}$ and covariance matrix $A^{-1}$ can be written as an iteration identical, except for an additive noise component, to the Gauss–Seidel iteration for solving the linear system $A\overline{x}=f$ to obtain $\overline{x}=A^{-1}f$; for more details see [goodman1989multigrid], as well as [Fox.C_Parker_2017_AcceleratedGibbsSampling]. Based on this observation, they proposed the so-called Multigrid Monte Carlo (MGMC) method—a random analogue of the multigrid method for solving discretised partial differential equations (PDEs). A careful analysis of the MGMC method is the focus of this paper. We assume that we have a positive definite operator $\mathcal{A}:H\to H$ on some function space $H$ on $\mathbb{R}^{d}$, for example $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ with positive integer $\alpha$ where $\Delta$ is the Laplace operator. A sufficiently accurate discretisation of $\mathcal{A}$ on some grid in $\mathbb{R}^{d}$ results in a possibly very large matrix $A\in\mathbb{R}^{n\times n}$. The goal of this paper is to show how to efficiently generate samples from $\mathcal{N}(\overline{x},A^{-1})$ for $n\rightarrow\infty$. We provide a rigorous theoretical justification for the optimal scaling of our method for large scale problems: the cost for generating an independent sample grows linearly with the problem size and the sampler is optimal in the continuum limit. Our analysis includes the important setting of a Gaussian random field conditioned on noisy data via a Bayesian approach; here, $\mathcal{A}$ is a finite-rank perturbation of some differential operator and the quantity of interest a functional of the Gaussian random field. Somewhat surprisingly, MGMC has not yet been considered in this context.
@@
    A grid-size-independent convergence theory for the MGMC method. The analysis shows that the first two moments, which fully characterise the Gaussian distribution, converge exponentially at a uniform rate to the target moments.
@@
    An extension of the MGMC algorithm, as well as its convergence theory, to the important situation of sampling Gaussian random fields conditioned on noisy data, i.e., to linear Bayesian inverse problems.
@@
    A corollary on the optimal complexity of MGMC to produce approximate samples whose mean and covariance converge to the mean and covariance of the limiting infinite-dimensional Gaussian random field.
@@
In contrast to stationary iterative methods on a single grid, such as the Gibbs sampler, the convergence rate of MGMC does not degenerate as the grid size tends to zero. The convergence analysis follows Fox and Parker [Fox.C_Parker_2017_AcceleratedGibbsSampling]. In that paper, they demonstrate that the convergence of the standard component-sweep Gibbs sampler is (in a suitable sense) equivalent to the deterministic (lexicographical) Gauss–Seidel iterative method, and that any matrix splitting used to generate a deterministic relaxation scheme also induces a stochastic relaxation that is a generalised Gibbs sampler; see also [norton2016tuning] where this analysis is extended also to non-Gaussian targets and to the popular Langevin and hybrid (or Hamiltonian) Monte Carlo samplers. Here, we extend the analysis in [Fox.C_Parker_2017_AcceleratedGibbsSampling] to MGMC. In doing so, we need to go beyond the derivation in [goodman1989multigrid] and beyond the standard multigrid theory (as presented for example in [Hackbusch.W_2016_IterativeSolutionLarge]) which provides convergence of the mean. Extra work is necessary to obtain the convergence of the covariance and to extend the analysis to the conditional case in a Bayesian setting. In particular, for linear functionals of the random field that are bounded in $H$, the crucial approximation and smoothing properties for the conditional case can be reduced to the corresponding properties in the unconditional case.
@@ 2.2 Linear Bayesian inverse problem
We are interested in generating samples from the posterior distribution $\mathcal{N}(\mu,\widetilde{\mathcal{A}}^{-1})$ and this paper focuses on the multigrid Monte Carlo sampler, which was introduced in [goodman1989multigrid] and is described in Section 3 below. The generated samples can be used to compute the expected value of a quantity of interest which is a functional $\mathcal{F}:H\rightarrow\mathbb{R}$. In the following we assume that $\mathcal{F}$ is a bounded linear functional of the form
@@
where $\chi$ is an element in $H$. In a slight abuse of notation, we will sometimes use $\mathcal{F}$ and its Riesz-representer $\chi$ interchangeably.
@@ 2.3 Discretisation
To realise such random variables, we follow [DaPrato.D_2014_book, Section 4.1.2] and let $(\varphi_{j})_{j\geq 1}$ be a complete $H$-orthonormal system (ONS) consisting of eigenfunctions of $\mathcal{A}^{-1}$ with corresponding eigenvalues $(1/\lambda_{j})_{j\geq 1}$. Then, $(\varphi_{j}/\sqrt{\lambda_{j}})_{j\geq 1}$ is a complete ONS of $V=D(\mathcal{A}^{1/2})\subset H$. For each $\chi\in H$, we define $(v,\chi)$ as the $L^{2}(\Omega)$-limit of the series
@@
the series in (2.20) converges the resulting random variable $(v,\chi)$ has the desired mean and covariance in (2.18)–(2.19). Moreover, following again [DaPrato.D_2014_book] the random field $v$ can be directly constructed on a suitably large Hilbert space without taking the coupling with $\chi\in H$.
@@ 4 Theory
This section presents our main theoretical results. We first put the relationship between random samplers and linear solvers already mentioned in Section 3.3 into a wider context: in Section 4.1 we argue that according to the pioneering work in [Fox.C_Parker_2017_AcceleratedGibbsSampling], the convergence of the Multigrid Monte Carlo sampler is closely related to the convergence of deterministic multigrid solvers for linear systems of equations. This is made more explicit in Section 4.2, where we show that the Multigrid Monte Carlo update Alg. 1 leaves the target distribution invariant, and Section 4.3, where we prove grid-independent convergence of the algorithm. Throughout the discussion we highlight the connection with the theory for deterministic multigrid methods; to streamline the presentation we defer proofs to the appendices. Finally, in Section 4.4, we apply and extend the theory to conditional distributions and Bayesian inverse problems introduced in Section 2.2. We conclude with a result on the optimality of the MGMG sampler for drawing samples from the target distribution in Section 4.5.
@@ 4.4 Application to linear Bayesian inverse problems
$0\leq x^{\top}A_{\ell}x+x^{\top}B_{\ell}\Gamma^{-1}B_{\ell}^{\top}x=x^{\top}\widetilde{A}_{\ell}x\leq x^{\top}\widetilde{A}_{\ell}x+x^{\top}L_{\ell}\left(D_{\ell}+B_{\ell}\Gamma^{-1}B_{\ell}^{\top}\right)^{-1}L_{\ell}^{\top}x=x^{\top}\widetilde{M}_{\ell}^{\text{(SGS)}}x\,,$
where the first inequality follows since the matrices $A_{\ell}$ and $\Gamma$ are assumed to be positive definite.
@@ 4.4.1 Approximation property for the perturbed problem
In the lemma above we assumed $\|M_{\ell}^{\text{(SGS)}}\|_{2}\leq C_{A}\|A_{\ell}\|_{2}$ for the unpurterbed precision matrix; this is a standard assumption (see e.g. [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.30]).
@@ 4.5 Optimality of MGMC
The MGMC update in Alg. 1 is efficient in the sense that is specified in Corollary 4.21 below. For this, we consider the Bayesian setting in Section 4.4 with a symmetric positive definite operator $\widetilde{\mathcal{A}}$ that has a compact inverse (which is not necessarily of trace class) and vectors $f_{L}$ and $F_{L}$ that are defined as in (2.10) and (2.16) respectively; we remind the reader that these two vectors are constructed with the help of the bounded linear operators $\mathcal{B}$ and $\mathcal{F}$ introduced in Section 2.2.
@@
Then, the Multigrid Monte Carlo algorithm is algorithmically optimal in the limit $n_{L}\rightarrow\infty$ in the sense that generating an (approximately) independent sample incurs a cost which grows no more than linearly in the number of unknowns. The generated samples are drawn from a distribution which is exponentially close to the target distribution, independent of the problem size. More specifically:
@@
    The cost of generating a new sample in the Markov chain with Alg. 1 is proportional to the number of unknowns $n_{L}$, for $n_{L}\to\infty$.
@@
Under Assumption 2.2, the sequence $(F_{L}^{\top}\theta_{L}^{*})_{L\in\mathbb{N}}$ with $\theta_{L}^{*}\sim\mathcal{N}(\widetilde{A}_{L}^{-1}f_{L},\widetilde{A}_{L}^{-1})$ converges in distribution to the $\mathbb{R}$-valued Gaussian random variable $(\widetilde{v},\mathcal{F})$ with mean and covariance given in Eqs. (2.21)–(2.22) (but with $\chi=\psi=\mathcal{F}$).
@@
Perhaps not very surprisingly, the notion of optimality of MGMC in Corollary 4.21 is closely related to that of multigrid solvers. For this, consider for example [Hackbusch.W_2016_IterativeSolutionLarge, Theorem 11.16] and [Hackbusch.W_2016_IterativeSolutionLarge, Remark 11.17] which shows that under certain conditions on the cycle parameter $\gamma$ the cost per multigrid cycle is bounded by a constant times the problem size $n_{L}$. Further, in [Hackbusch.W_2016_IterativeSolutionLarge, Section 11.5.5] the application of a nested multigrid iteration to PDE based problems is considered. It is shown there that this approach can reduce the error to the order of the discretisation error at a cost that can be bounded by a constant times $n_{L}$.